6. Brewers’ Company's (BRW) common stock is currently trading for $25.00 per share. The stock is expected to pay a $2.50 dividend at the end of the year and BRW's equity cost of capital rE is 14%. (5pts) All handwritten no financial calculator.

(a)If the dividend payout rate (of 75%) is expected to remain constant, then what is the expected growth rate in BRW's earnings? (12pts)

(b)Suppose that BRW has a new investment opportunity that is expected to yield a return of 12%, should BRW reduce its dividend payout rate to 70% in order to invest in this project? Calculate and explain your reasoning. (3 pts)

(c) Could you have arrived at your answer to (b) without calculations? Explain.  
Answer #1

P = 01/(R-8) g=R-D1/P .14 - 2.5/25 4.00% b) Now the earning must be EPS = 2.5/.75 3.33 With a change in dividend payout to 70